OpenStudy (anonymous):

I got part A right, but the rest…idk. Solve the problems below that pertain to the profit from the sale of items. (a) Marketing tells you that if you set the price of an item at $10 then you will be unable to sell it, but that you can sell 560 items for each dollar below $10 that you set the price. Suppose your fixed costs total $4000, and your marginal cost is $2 per item. What is the most profit you can make? _4960_ $ . when the price is $_6_ . . (b) How are your answers to the problem affected if the cost per item for the x items, instead of being simply $2, decreases below $2 in pr

OpenStudy (anonymous):

well if you lower the cost of making something what would happen to your profit

OpenStudy (anonymous):

I know it's basically the same as part A with just a tiny variation of the Cost function

OpenStudy (anonymous):

But I need actual numbers. obviously the profit would increase

OpenStudy (anonymous):

it's asking for an answer but it doesn't give you how low the cost per item decreases?

OpenStudy (anonymous):

Wait it didn't post the whole question! Grrrr

OpenStudy (anonymous):

Hey wait you're an engineer?! What kind?

OpenStudy (anonymous):

not an engineer i'm still in school, Electrical and Computer

OpenStudy (anonymous):

What year?

OpenStudy (anonymous):

Im a freshman undecided Engineer

OpenStudy (anonymous):

2.5 to 3 ish

OpenStudy (anonymous):

nice! yeah I know it's a weird time format because people do things different sometimes. that's cool :) Okay okay Calculus. I need help with the section that talks about "decreases below $2 in proportion to x by 1 cent for each 28 items produced"

OpenStudy (anonymous):

anyways your proportion is \[.01(\frac{x}{28}\]

OpenStudy (anonymous):

setting up that part of the equation is my downfall

OpenStudy (anonymous):

i think but i'm not sure how they want to deal with fractional numbers -.-

OpenStudy (anonymous):

right but does that go right back into the Cost function?

OpenStudy (anonymous):

um yes it goes into the maginal part

OpenStudy (anonymous):

so you have \[-(2x-.01(\frac{x}{28}))\]

OpenStudy (anonymous):

i think your marginal part was -2x correct?

OpenStudy (anonymous):

so C=4000 +( 2x-.01(x/28))

OpenStudy (anonymous):

it'd help to see your actual cost function

OpenStudy (anonymous):

My original Cost was C=4000+2x nothing fancy

OpenStudy (anonymous):

yes so it'd be that

OpenStudy (anonymous):

i was thining of a profit function

OpenStudy (anonymous):

in which your costs would be negative

OpenStudy (anonymous):

hmm but I think I did that and I got about the same Profit as before...which wouldn't make sense

OpenStudy (anonymous):

yeah the Profit function was P= 10x-(x^2 / 560) - (4000+2x) for part A

OpenStudy (anonymous):

well you're only subtracting a cent per 28 items

OpenStudy (anonymous):

so it'd make sense that your answer is pretty close to your original answer

OpenStudy (anonymous):

hmm then maybe it doesn't want decimals?

OpenStudy (anonymous):

Okay then, I can do that one later, but can you look at one more problem? or are you as done as I am?

OpenStudy (anonymous):